# A High-Quality Dataset and Reliable Evaluation for Interleaved Image-Text Generation

### v1

@@ 6 Conclusion
In this paper, we present InterSyn, a large-scale, high-quality multimodal dataset designed for instruction-following and interleaved image-text generation. Constructed via the fully automated SEIR method, InterSyn combines scale, diversity, and fidelity, supporting multi-turn dialogues where each response is refined to achieve not only semantic completeness but also tight image-text synergy—ensuring that visual and textual modalities complement each other to convey meaning collaboratively. To complement InterSyn, we introduce SynJudge, a multi-dimensional automatic evaluator specifically designed to assess interleaved outputs across four key dimensions, including a dedicated metric for image-text synergy. Unlike traditional metrics focused on surface-level alignment or consistency, SynJudge emphasizes the semantic interplay between images and text, rewarding complementary relationships while penalizing redundancy or disjointness. Extensive experiments validate the effectiveness of both InterSyn and SEIR. Models fine-tuned on InterSyn consistently outperform strong baselines, showing notable improvements in instruction alignment, multimodal reasoning, and especially the ability to produce coherent, synergistic interleaved content. We believe this work lays a solid foundation for future research in scalable multimodal data generation, robust synergy-centric evaluation, and the development of general-purpose multimodal intelligence systems that understand and communicate across modalities in a truly integrated manner.
Table 4: Generator performance evaluated by human judge and SynJudge.
Generator Dimension Human SynJudge
TCC ICC IQ ITS TCC ICC IQ ITS
Anole mean 2.86 2.75 2.69 2.05 2.89 2.81 2.72 2.06
variance 1.47 2.05 1.87 2.47 1.47 2.11 1.93 2.55
DDiT mean 0.38 3.51 3.29 0.37 0.28 3.67 3.34 0.26
variance 1.17 0.83 0.85 1.21 0.86 0.75 0.92 0.87
Emu3 mean 3.38 3.86 3.87 3.26 3.37 3.92 3.85 3.37
variance 0.86 0.86 0.59 1.31 0.74 0.66 0.80 1.16
VILA-U mean 2.25 3.42 3.15 2.11 2.26 3.52 3.17 1.99
variance 2.58 0.91 0.97 3.01 2.73 0.92 1.12 2.96
Liquid mean 2.88 3.82 3.67 3.00 2.87 3.86 3.76 3.06
variance 0.76 0.69 0.70 1.42 0.71 0.75 0.85 1.36
Janus-Pro mean 2.93 3.25 3.16 2.55 2.96 3.30 3.11 2.62
variance 0.95 1.21 0.99 1.75 0.96 1.09 1.11 1.87
Show-o mean 3.32 3.65 3.52 3.10 3.49 3.79 3.57 3.30
variance 1.31 0.94 0.89 1.94 1.12 0.82 0.96 1.70
Show-o-Turbo mean 3.48 3.77 3.53 3.33 3.50 3.89 3.64 3.45
variance 1.12 0.88 0.94 1.60 1.06 0.86 1.01 1.44
VARGPT mean 2.60 0.94 0.94 0.55 2.55 0.94 0.87 0.67
variance 0.66 2.37 2.32 1.99 0.30 2.39 2.17 2.16
Gemini+Flux mean 3.94 4.06 4.43 3.81 3.97 4.12 4.48 3.84
variance 0.94 0.58 0.47 0.90 0.57 0.71 0.64 1.11
GPT-4o+DALL-E mean 4.05 4.08 4.41 3.94 3.99 4.10 4.45 3.87
variance 0.37 0.48 0.57 0.64 0.65 0.81 0.58 1.16
InterSyn mean 4.41 4.46 4.47 4.51 4.39 4.49 4.44 4.53
variance 0.55 0.55 0.53 0.57 0.61 0.63 0.45 0.51
((a)) Performance of Anole.
AR IR TCC ICC IQ ITS
null null 3.09 3.01 2.92 2.26
0 0 3.33 3.17 2.92 2.77
1 1 3.37 3.21 3.07 2.71
2 2 3.41 3.27 3.03 2.79
3 0 3.47 3.20 3.03 2.85
3 1 3.51 3.25 3.11 2.93
3 2 3.49 3.3 3.08 2.91
3 3 3.52 3.32 3.1 2.94
((b)) Performance of VILA-U.
AR IR TCC ICC IQ ITS
null null 2.46 3.72 3.37 2.19
0 0 3.07 3.83 3.37 3.13
1 0 3.1 3.8 3.34 3.2
2 0 3.17 3.79 3.39 3.27
3 0 3.17 3.81 3.36 3.27
3 1 3.17 3.81 3.4 3.3
3 2 3.21 3.85 3.37 3.29
3 3 3.19 3.83 3.39 3.33
Appendix A Appendix overview
@@
  • Comparison of datasets and Samples of InterSyn are shown in Sec. B.
  • Supplementary analysis of experimental Data are shown in Sec. C.
  • Question templates and topic hierarchy are shown in Sec. D.
  • The evaluation dimensions for question and answer are shown in Sec. E.
  • Prompts used in this work are shown in Sec. F.
  • The benchmark samples are shown in Sec. G.
  • Human annotation platform are shown in Sec. H.
  • Limitations of this study are shown in Sec. I.
Appendix B Comparison of Datasets and InterSyn Samples
B.1 Comparison of Datasets
The table provides a comprehensive comparison of InterSyn with representative multimodal datasets and benchmarks. Existing datasets such as MMC4 and OBELICS primarily rely on large-scale web-crawled corpora, often lacking instruction-following capabilities and multi-turn structures. Other resources like CoMM and ShareGPT4V improve data cleanliness but remain limited to single-turn interactions without tight semantic supervision.
@@ B.1 Comparison of Datasets
In contrast, InterSyn is the first to offer a large-scale, multi-turn, instruction-following dataset specifically designed for interleaved image-text generation. Built with the SEIR method, InterSyn not only ensures high-quality visual-textual synergy but also scales to 1.8 million samples—orders of magnitude larger than existing benchmarks. Its emphasis on dialogue coherence, iterative refinement, and synergistic multimodal responses fills a critical gap in current resources, laying the groundwork for developing and evaluating truly unified multimodal generation models.
@@
B.2 Single-Turn Samples
@@
B.3 Multi-Turn Samples
@@
Appendix C Supplementary Analysis of Experimental
C.1 Symbols and Notations for SEIR Method
@@
  • $\mathcal{H}^{(t-1)}$: History of the conversation up to turn $t-1$, represented as $\{(q^{(i)},a^{(i)},i^{(i)})\}_{i=1}^{t-1}$.
@@
C.2 Evaluation Metrics for Judges and Generators
@@
Here, $x_{i,d}^{M}$ and $x_{i,d}^{H}$ denote the scores assigned by the model-based and human judges respectively.
RMSE quantifies the deviation between a model-based judge’s scores and those of human judge in dimension $d$. Lower RMSE values indicate higher alignment with human preferences, and thus higher reliability of the model-based judge.
This evaluation framework provides a comprehensive analysis of the performance of the 11 generators across multiple dimensions, ensuring objective comparison from both human and model-based perspectives.
C.3 Models Used in the SEIR Method
To construct the dataset, we adopt Qwen2.5-32B-Instruct as the language model ($\mathcal{M}_{L}$), InternVL2_5-26B as the vision-language model ($\mathcal{M}_{V}$), and FLUX.1-dev as the text-to-image generation model ($\mathcal{M}_{G}$). These open-source models are used as the default configuration in our SEIR framework. Importantly, the framework is modular by design—each component can be substituted with other models, offering flexibility for different deployment environments or research needs.
@@ C.3 Models Used in the SEIR Method
C.4 Hyperparameters Used for Training Judge Model
We fine-tuned two large multimodal models, InternVL2.5-8B and QwenVL2.5-7B, as judge models to evaluate the interleavde image-text response quality. We followed common practices for large-scale model fine-tuning, applying weight decay regularization, learning rate warmup, and gradient clipping to ensure training stability. All experiments were conducted using mixed-precision training on distributed GPU clusters.
@@ C.4 Hyperparameters Used for Training Judge Model
C.5 Detailed Analysis of the Judge’s Scoring
@@
Judge Accuracy Comparison.
Table LABEL:tab:big-data-ratio reports the evaluation accuracy of different judges across models and dimensions. Considering the inherent subjectivity in evaluation, we define a deviation within 1 point between model and human scores as an acceptable margin of subjective error. Accordingly, the reported accuracy reflects the proportion of samples where the model’s score differs from human judgment by no more than one point. InternVL_trained also shows strong performance with an average accuracy of 94.5%, slightly lower but still significantly better than non-fine-tuned baselines. In contrast, GPT-4o and InternVL base models exhibit noticeably lower accuracy, around 86.5%, indicating that although they are strong in generation, they are less reliable for fine-grained evaluation without specialized tuning. Based on these findings, we select QwenVL_trained as the backbone model for constructing SynJudge, ensuring both accuracy and consistency in downstream evaluations.
@@
Table 13: Evaluation accuracy comparison across judges
@@
Average all 0.865 0.866 0.945 0.875 0.954
C.6 Detailed analysis of the capabilities of different Generators
@@
Figure 7 presents the mean and variance of evaluation scores across TCC, ICC, IQ, and ITS for different generators. Several important observations emerge from the results. First, DDiT exhibits the lowest mean scores in both TCC (0.38) and ITS (0.37), indicating poor content coverage and weak image-text synergy. Second, VARGPT shows mean scores below 1 across ICC, IQ, and ITS, suggesting significant deficiencies in visual generation capabilities and multimodal alignment. Third, VILA-U demonstrates the highest variance in TCC and ITS among all models, implying that its performance is highly unstable across different questions. In contrast, Gemini+Flux and GPT+DALL-E achieve mean scores above 4.0 across all evaluation dimensions, reflecting generally strong performance. However, their relatively high variance in ITS reveals that they still struggle with maintaining image-text consistency and complementarity across samples. Most notably, InterSyn consistently outperforms all other generators across all four dimensions, achieving the highest mean scores while maintaining the lowest variance. This indicates not only superior quality but also high stability and robustness in both textual and visual generation. These findings collectively highlight the importance of both quality and consistency for robust multimodal generation, and demonstrate the synergy of SEIR in constructing high-quality, stable datasets.
@@
Appendix D Question Templates and Topic Hierarchy
D.1 Question Templates
@@
Can you draw a picture of ***? Besides, can you give me some popular science knowledge about ***?
@@
What kind of wonders can be seen in ***? Im really curious about what that scene will be like. Can you show it to me?
@@
Hey, can you tell me a legend about *** and draw a vivid picture?
D.2 Topic Hierarchy
@@
Appendix E Evaluation Dimensions for Dataset Quality
E.1 Question Evaluation Dimensions
@@
Reasonableness of Expression: The question statement is smooth, without any grammatical errors, and the words are used accurately and appropriately. For example, "Please introduce the Great Wall to me and also give me a picture of the Great Wall" is a reasonable expression; while "Tell me about the Great Wall, and give me a picture" has a problem of confused expression. Such questions will affect the model’s understanding of the intention and make it difficult to give an accurate answer.
@@
Clarity of Requirements: Clearly indicate that the model is required to provide both text and image responses simultaneously. For instance, "Introduce the appearance characteristics of Notre-Dame de Paris and provide a high-definition frontal picture", which clearly puts forward the dual requirements of text description and image acquisition; if the question is just "What does Notre-Dame de Paris look like", without clearly stating the image requirement, it does not meet the requirements and cannot effectively guide the model to give a comprehensive response.
@@
Focus of the Theme: The question revolves around a single and clear theme and will not jump between multiple unrelated themes. For example, "Introduce the geographical features of Mount Fuji and attach a distant view of Mount Fuji", with the theme focused on Mount Fuji; while "Tell me about Mount Fuji and then talk about the Eiffel Tower, and give two corresponding pictures", which involves two different themes, may lead to unclear logical answers from the model and is not conducive to the standardized construction of the dataset.
@@
Feasibility and Clarity: Based on common sense judgment, the content involved in the question is something that the model has the ability to answer through language and images, and there is no way of multiple interpretations, and the model can accurately grasp the questioner’s intention. For example, "Describe the living habits of giant pandas and give a picture of a panda eating bamboo", the model can answer based on its existing knowledge reserve and image generation ability, and the intention is clear; however, "Tell me what it’s like for a person to take a bath in volcanic magma and give a picture", such questions seriously deviate from reality and lack scientific basis. The model can neither answer based on existing knowledge, nor is there a real-world reference for image generation, which will lead to absurd and meaningless generation results and greatly reduce the reliability and practicality of the dataset.
@@
Appropriateness of Length: The length of the question is moderate, which not only contains enough key information to guide the model to generate high-quality answers but also is not too long and complicated for the model to grasp the key points. Generally speaking, short and concise questions are helpful for the model to quickly understand the intention, such as "Introduce the Forbidden City and give a panoramic picture"; but being too short may lack sufficient information, such as "Forbidden City, picture"; and overly long and cumbersome questions, like "Please introduce in detail the process of the changes of the Forbidden City since its construction in the Ming Dynasty through various dynasties, including the evolution of architectural styles, the transformation of functional uses and other aspects, and provide a high-definition panoramic picture that can comprehensively display the current overall layout of the Forbidden City. At the same time, ensure that the picture contains the main palaces, courtyards, city walls and other iconic elements of the Forbidden City", may cause confusion when the model processes it. The ideal length can be determined according to practical experience and testing. Usually, about 15 - 50 words is more appropriate, which can convey the requirements completely and also facilitate the model to process efficiently.
E.2 Interleaved Image-Text Answer Evaluation Dimensions

### v2

@@ 5 Conclusion
In this paper, we present InterSyn, a large scale, high quality multimodal dataset designed for instruction-following and interleaved image-text generation. Constructed via the fully automated SEIR method, InterSyn combines scale, diversity, and fidelity, supporting multi-turn dialogues where each response is refined to achieve not only semantic completeness but also tight image-text synergy—ensuring that visual and textual modalities complement each other to convey meaning collaboratively. To complement InterSyn, we introduce SynJudge, a multi-dimensional automatic evaluator specifically designed to assess interleaved outputs across four key dimensions, including a dedicated metric for image-text synergy. Unlike traditional metrics focused on surface-level alignment or consistency, SynJudge emphasizes the semantic interplay between images and text, rewarding complementary relationships while penalizing redundancy or disjointness. Extensive experiments validate the effectiveness of both InterSyn and SEIR. Models fine-tuned on InterSyn consistently outperform strong baselines, showing notable improvements in instruction alignment, multimodal reasoning, and especially the ability to produce coherent, synergistic interleaved content. We believe this work lays a solid foundation for future research in scalable multimodal data generation, robust synergy-centric evaluation, and the development of general-purpose multimodal intelligence systems that understand and communicate across modalities in a truly integrated manner.
Reproducibility Statement. To facilitate the verification of our findings and to support future research, we are committed to making our work fully reproducible. The complete codebase, including scripts for the SEIR data generation pipeline, SynJudge training, and all evaluation protocols, will be made publicly available on GitHub upon publication. We will also release the full InterSyn dataset (1.8M samples), all benchmark sets used for our main experiments, and the trained weights of our final SynJudge evaluator. The core of our methodology relies on publicly available models, and all prompts, model configurations, and hyperparameters are extensively detailed in the appendix to ensure that our results can be precisely replicated.
Appendix A Statement on LLMs Usage
The authors used large language models (LLMs) during the writing process solely for language refinement and editing. It should be explicitly stated that LLMs were not employed in any core aspects of the research, including the formulation of research ideas, the design of methodologies, the execution of experiments, or the development of conclusions. All scholarly contributions were made independently by the authors.
Appendix B Appendix overview
@@
  • Comparison of datasets and Samples of InterSyn are shown in §C.
  • Supplementary analysis of experimental Data are shown in §D.
  • Question templates and topic hierarchy are shown in §E.
  • The evaluation dimensions for question and answer are shown in §F.
  • Prompts used in this work are shown in §G.
  • The benchmark samples are shown in §H.
  • Human annotation platform are shown in §I.
  • Limitations of this study are shown in §J.
Appendix C Comparison of Datasets and InterSyn Samples
C.1 Comparison of Datasets
The table provides a comprehensive comparison of InterSyn with representative multimodal datasets and benchmarks. Existing datasets such as MMC4 and OBELICS primarily rely on large scale web-crawled corpora, often lacking instruction-following capabilities and multi-turn structures. Other resources like CoMM and ShareGPT4V improve data cleanliness but remain limited to single-turn interactions without tight semantic supervision.
@@ C.1 Comparison of Datasets
In contrast, InterSyn is the first to offer a large scale, multi-turn, instruction-following dataset specifically designed for interleaved image-text generation. Built with the SEIR method, InterSyn not only ensures high quality visual-textual synergy but also scales to 1.8 million samples—orders of magnitude larger than existing benchmarks. Its emphasis on dialogue coherence, iterative refinement, and synergistic multimodal responses fills a critical gap in current resources, laying the groundwork for developing and evaluating truly unified multimodal generation models.
@@
C.2 Single-Turn Samples
@@
C.3 Multi-Turn Samples
@@
Appendix D Supplementary Analysis of Experimental
D.1 Symbols and Notations for SEIR Method
@@
  • $\mathcal{H}^{(t-1)}$: History of the conversation up to turn t1t-1, represented as $\{(q^{(i)},a^{(i)},i^{(i)})\}_{i=1}^{t-1}$.
@@
D.2 Evaluation Metrics for Judges and Generators
@@
Here, $x_{i,d}^{M}$ and $x_{i,d}^{H}$ denote the scores assigned by the model-based and human judge respectively. RMSE quantifies the deviation between a model-based judge’s scores and those of human judge in dimension $d$. Lower RMSE values indicate higher alignment with human preferences, and thus higher reliability of the model-based judge.
[…]
$\text{A@1}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(|x_{i,d}^{M}-x_{i,d}^{H}|\leq 1)$ (8)
where $\mathbb{I}(\cdot)$ is the indicator function, which is 1 if the condition is true and 0 otherwise. A higher A@1 score indicates that the judge’s scores are more frequently in close agreement with human evaluators, reflecting greater practical reliability.
This evaluation framework provides a comprehensive analysis of the performance of all the generators across multiple dimensions, ensuring objective comparison from both human and model-based perspectives.
D.3 Models Used in the SEIR Method
To construct the dataset, we adopt Qwen2.5-32B-Instruct as the language model ($\mathcal{M}_{L}$), InternVL2_5-26B as the vision-language model ($\mathcal{M}_{V}$), and FLUX.1-dev as the text-to-image generation model ($\mathcal{M}_{G}$). These open-source models are used as the default configuration for our SEIR framework. Importantly, the framework is modular by design—each component can be substituted with other models, offering flexibility for different deployment environments or research needs.
@@ D.3 Models Used in the SEIR Method
D.4 Hyperparameters Used for Training Judge Model
We fine-tuned two large multimodal models, InternVL2.5-8B and QwenVL2.5-7B, as judge models to evaluate the interleaved image-text content completeness. We followed common practices for large scale model fine-tuning, applying weight decay regularization, learning rate warmup, and gradient clipping to ensure training stability. All experiments were conducted using mixed-precision training on distributed GPU clusters.
@@ D.4 Hyperparameters Used for Training Judge Model
D.5 Further verification of SEIR effectiveness
((a)) Performance of Anole.
AR IR TCC ICC IQ ITS
null null 3.09 3.01 2.92 2.26
0 0 3.33 3.17 2.92 2.77
1 1 3.37 3.21 3.07 2.71
2 2 3.41 3.27 3.03 2.79
3 0 3.47 3.20 3.03 2.85
3 1 3.51 3.25 3.11 2.93
3 2 3.49 3.3 3.08 2.91
3 3 3.52 3.32 3.1 2.94
((b)) Performance of VILA-U.
AR IR TCC ICC IQ ITS
null null 2.46 3.72 3.37 2.19
0 0 3.07 3.83 3.37 3.13
1 0 3.1 3.8 3.34 3.2
2 0 3.17 3.79 3.39 3.27
3 0 3.17 3.81 3.36 3.27
3 1 3.17 3.81 3.4 3.3
3 2 3.21 3.85 3.37 3.29
3 3 3.19 3.83 3.39 3.33
D.6 SynJudge Train/Test Data
Our full annotated data contains 48,000 samples. These 48,000 questions were generated by the SEIR method through three iterations, with the question template and topic hierarchy ensuring sufficient diversity across topics and conversational query style.
To obtain a broad distribution of multimodal question–answer (QA) outputs, these 48,000 questions were randomly assigned to different generators, which then produced responses. This strategy ensures that the resulting QA pairs cover a wide quality spectrum. After generation, all question–answer pairs were annotated by trained human annotators for TCC, ICC, IQ, and ITS over a two-week period.
The annotated dataset was split into a training set (80%) and a test set (20%). The training set was used to fine-tune candidate judge models (e.g., QwenVL_trained, InternVL_trained), while the test set was reserved for evaluation. For the evaluation, we used RMSE as the metric to measure how closely the scores from different judges align with ground-truth human annotations: a lower RMSE indicates a better judge. This evaluation process is what leads to the selection of our final model, SynJudge.
D.7 Detailed Analysis of the Judge’s Scoring
@@
Judge Agreement Analysis.
Table LABEL:tab:big-data-ratio provides a detailed breakdown of the Human Agreement within Tolerance (A@1) for each judge across all evaluated generators and dimensions. As formally defined in Appendix §D.2, this metric reflects the proportion of scores where the absolute difference between the model and human judge is no more than one point ($\tau=1$), which we consider an acceptable margin for subjective tasks.
The results reinforce our findings from the main paper. The finetuned judges consistently achieve higher agreement rates. QwenVL_trained achieves the highest average A@1 of 95.4%. InternVL_trained also shows strong performance at 94.5%, significantly better than the non-finetuned baselines. In contrast, zero-shot models like GPT-4o and QwenVL exhibit noticeably lower agreement, around 86.5% and 87.5% respectively, indicating that they are less reliable for fine-grained evaluation without specialized tuning. These findings further justify our selection of QwenVL_trained as the backbone for SynJudge.
@@
Table 15: Evaluation accuracy A@1 comparison across judges
@@
Average A@1 0.865 0.866 0.945 0.875 0.954
D.8 Detailed analysis of the capabilities of different Generators
@@
Figure 7 presents the mean and variance of evaluation scores across TCC, ICC, IQ, and ITS for different generators. Several important observations emerge from the results. First, DDiT exhibits the lowest mean scores in both TCC (0.38) and ITS (0.37), indicating poor content coverage and weak image-text synergy. Second, VARGPT shows mean scores below 1 across ICC, IQ, and ITS, suggesting significant deficiencies in visual generation capabilities and multimodal alignment. Third, VILA-U demonstrates the highest variance in TCC and ITS among all models, implying that its performance is highly unstable across different questions. In contrast, Gemini+Flux and GPT-4o+DALL-E achieve mean scores above 4.0 across all evaluation dimensions, reflecting generally strong performance. However, their relatively high variance in ITS reveals that they still struggle with maintaining image-text consistency and complementarity across samples. Most notably, SEIR method consistently outperforms all other generators across all four dimensions, achieving the highest mean scores while maintaining the lowest variance. This indicates not only superior quality but also high stability and robustness in both textual and visual generation. These findings collectively highlight the importance of both quality and consistency for robust multimodal generation, and demonstrate the synergy of SEIR in constructing high quality, stable datasets.
@@
Appendix E Question Templates and Topic Hierarchy
In our framework, question templates are designed to model the style of human queries rather than to encode domain knowledge. Specifically, these templates capture the recurrent syntactic and pragmatic structures through which users naturally formulate requests. For example, variations such as “Please recommend some equipment needed for hiking,”, “Tell me some equipment needed for hiking,” or “Could you tell me what equipment is needed for hiking?” all convey the same underlying intent but differ in their linguistic form. By systematically enumerating such structures, we obtain broad coverage of syntactic patterns for conversational queries (e.g., “can you …,” “please …”), while leaving the semantic content to be drawn from a large topic hierarchy. This separation ensures that the linguistic diversity of user queries can be effectively modeled: the question template specifies how a request is asked, whereas the the 3500-topic hierarchy defines what the request is about. The combination of these two components enables our dataset to achieve high diversity while faithfully reflecting real-world user interaction styles.
E.1 Question Templates
@@
Can you draw a picture of ***? Besides, can you give me some science knowledge about ***?
@@
What kind of wonders can be seen in ***? Im really curious about the scene. Can you show it to me?
@@
Please introduce *** and draw about it.
Hey, can you tell me *** and draw a picture?
E.2 Topic Hierarchy
To provide a clearer picture of the dataset’s composition, the table 16 below shows the distribution across our 8 primary domains, including the number of categories and fine-grained topics within each. As you can see, the data is well-distributed across diverse topics like Natural Scenery (19.88%), Cultural Scenery (15.87%), and Animals (14.77%), ensuring broad and deep coverage.
Table 16: Data distribution of InterSyn across 8 primary domains. The table details the percentage domain distribution (Domain Dist.), distribution per Category (Dist. per Cat.), number of categories (# Categories), and number of topics for each domain (# Topics), demonstrating broad and deep coverage.
Domain Domain Dist. (%) Dist. per Cat. (%) # Categories # Topics
Animals 14.77 1.64 9 517
Plants 10.13 1.27 8 355
Natural Scenery 19.88 2.48 8 695
Cultural Scenery 15.87 1.98 8 555
Objects 10.65 1.33 8 373
Activities 11.20 1.40 8 392
Food 6.71 0.84 8 235
Culture 10.80 1.35 8 378
@@ E.2 Topic Hierarchy
Appendix F Evaluation Dimensions for Dataset Quality
F.1 Question Evaluation Dimensions
@@
Reasonableness of Expression: The question statement is smooth, without any grammatical errors, and the words are used accurately and appropriately. For example, ”Please introduce the Great Wall to me and also give me a picture of the Great Wall” is a reasonable expression; while ”Tell me about the Great Wall, and give me a picture” has a problem of confused expression. Such questions will affect the model’s understanding of the intention and make it difficult to give an accurate answer.
@@
Clarity of Requirements: Clearly indicate that the model is required to provide both text and image responses simultaneously. For instance, ”Introduce the appearance characteristics of Notre-Dame de Paris and provide a high-definition frontal picture”, which clearly puts forward the dual requirements of text description and image acquisition; if the question is just ”What does Notre-Dame de Paris look like”, without clearly stating the image requirement, it does not meet the requirements and cannot effectively guide the model to give a comprehensive response.
@@
Focus of the Theme: The question revolves around a single and clear theme and will not jump between multiple unrelated themes. For example, ”Introduce the geographical features of Mount Fuji and attach a distant view of Mount Fuji”, with the theme focused on Mount Fuji; while ”Tell me about Mount Fuji and then talk about the Eiffel Tower, and give two corresponding pictures”, which involves two different themes, may lead to unclear logical answers from the model and is not conducive to the standardized construction of the dataset.
@@
Feasibility and Clarity: Based on common sense judgment, the content involved in the question is something that the model has the ability to answer through language and images, and there is no way of multiple interpretations, and the model can accurately grasp the questioner’s intention. For example, ”Describe the living habits of giant pandas and give a picture of a panda eating bamboo”, the model can answer based on its existing knowledge reserve and image generation ability, and the intention is clear; however, ”Tell me what it’s like for a person to take a bath in volcanic magma and give a picture”, such questions seriously deviate from reality and lack scientific basis. The model can neither answer based on existing knowledge, nor is there a real-world reference for image generation, which will lead to absurd and meaningless generation results and greatly reduce the reliability and practicality of the dataset.
@@
Appropriateness of Length: The length of the question is moderate, which not only contains enough key information to guide the model to generate high quality answers but also is not too long and complicated for the model to grasp the key points. Generally speaking, short and concise questions are helpful for the model to quickly understand the intention, such as ”Introduce the Forbidden City and give a panoramic picture”; but being too short may lack sufficient information, such as ”Forbidden City, picture”; and overly long and cumbersome questions, like ”Please introduce in detail the process of the changes of the Forbidden City since its construction in the Ming Dynasty through various dynasties, including the evolution of architectural styles, the transformation of functional uses and other aspects, and provide a high-definition panoramic picture that can comprehensively display the current overall layout of the Forbidden City. At the same time, ensure that the picture contains the main palaces, courtyards, city walls and other iconic elements of the Forbidden City”, may cause confusion when the model processes it. The ideal length can be determined according to practical experience and testing. Usually, about 15 - 50 words is more appropriate, which can convey the requirements completely and also facilitate the model to process efficiently.
F.2 Interleaved Image-Text Answer Evaluation Dimensions